OpenStudy (anonymous):

From an environmental study of a certain community it is estimated that there will be Q(p)=p^2 + 3p +1200 units of a harmful pollutant in the air when the population is p thousand. The population is currently 30,000 people and is increasing at the rate of 2000 people per year. At what rate is the level of air pollution increasing currently?

OpenStudy (anonymous):

can you do derivatives?

OpenStudy (anonymous):

you need to find the rate of change of Q with respect to p... differentiate the expression for Q(p) with respect to p.

OpenStudy (anonymous):

Yes derivatives. I'm just really confused about the derivatives

OpenStudy (anonymous):

when you take the derivative of a term like x^2, you reduce its exponent by 1, and you multiply it by whatever the exponent used to be... so (d/dx) of x^2 becomes 2x

OpenStudy (anonymous):

differentiate Q(p) with respect to p: Q(p)=p^2 + 3p +1200 (d/dp) Q(p) = 2p + 3 <<<--- can you see why?

OpenStudy (anonymous):

I put dQ/dt= 2000(2p + 3) =4000p + 6000 =4000(30,000) + 6000 =120006000

OpenStudy (anonymous):

I think once you have dQ/dp = 2p + 3 then you substitute in p = 30,000... that's the current population. the rate of change of Q when p = 30,000 will be 2(30000) + 3 = 60,003 the population increase of 2000 is a "trick" to distract you... you just need to worry about the rate of change of the pollution, not the population

OpenStudy (anonymous):

I think I made an unpleasant error... I am very sorry!

OpenStudy (anonymous):

that expression I just showed was the dQ/dp... but you were correct when you were thinking that you needed dQ/dt, not dQ/dp

OpenStudy (anonymous):

uggh! Apologies again... that was an awful error I made.

OpenStudy (anonymous):

thats ok! I appreciate your help! so would the expression and answer be the same?

OpenStudy (anonymous):

I am not sure... I was trying to look it up somewhere and learn it (again) really quick.

OpenStudy (anonymous):

I'm thinking maybe you might want to copy and paste your question into a new question and close this one... the discussion above is confusing... might be better to start fresh and get someone who knows for sure how to do this.

OpenStudy (anonymous):

Ok thank u!